# Semi-analytical covariance matrices for two-point correlation function for DESI 2024 data

### v2

@@ 2.2 Comparison measures for covariance matrices
We employ the same three compact measures of covariance matrix similarity as in RascalC validation for early DESI data [31]. We use RascalC precision (inverse covariance) matrix and the mock sample covariance matrix because the latter is more noisy and thus less stable to inversion. We also find this ordering more interpretable: some additional properties do not hold with the other one, as noted below.
@@
(2.9) which is used to optimize shot-noise rescaling, and related to the log-likelihood of the sample covariance under the assumption that the RascalC covariance is truly describing the distribution of mock clustering measurements555This log-likelihood relation does not hold if the KL divergence is computed between sample precision and theoretical covariance matrices. [26].
@@
(2.10) -
@@
(2.11) which is distributed like a reduced with degrees of freedom under the assumption that RascalC covariance describes the distribution of the mock clustering measurements666Again, if we computed the mean reduced between the mock precision and theoretical covariance matrices, it would not follow the reduced distribution exactly..
The first two are sensitive to deviations in different ways, but their expectation values would not be zero even if RascalC covariance matrices perfectly matched the true underlying covariance. The last measure, mean reduced , is more sensitive to the overall scaling between the covariance matrices, while deviations in different directions can cancel each other.
The expectation values and standard deviations for all the measures have been derived in [31]; the result for is exact while the other two are approximate. We have performed empirical validations using 10,000 samples with a unit covariance matrix and the desired number of bins. We have found significant deviations for KL divergence in observable space (due to more multipoles and thus total bins considered in this paper) and smaller ones for in parameter space; in these cases, we reverted to the empirical means and standard deviations.
@@ 2.3 Fisher projection to the space of model parameters
As in [31], we approximate the covariance matrices in the parameter space through the inverse of the Fisher matrix. For RascalC results, we a simple expression without inversion bias corrections (which are small for our semi-analytical covariance matrices):
is the observable-space RascalC covariance matrix cut to the fit range for the respective model encompassing observables, and is the Jacobian of the model, specifically the matrix of derivatives of the vector of radially-binned 2PCF multipoles with respect to the parameter vector :
@@ 3 Covariance for projected Legendre moments of 2PCF revisited
Because of the great advantage of using few estimators and hence a smaller covariance matrix, multipole moments are a compression favorable to many angular bins. A past RascalC extension introduced the estimators for their covariance, assuming weighting by Legendre polynomials during pair counting, or infinitesimally narrow angular bins [29]. At the same time, the 2-point correlation function estimation library widely used in DESI, pycorr777https://github.com/cosmodesi/pycorr, estimates the radially binned Legendre moments of the 2-point correlation function from the radially and angularly binned estimates (Eq. 2.1) with angular bins (after wrapping):
are the projection factors, which do not depend on radial bins or the tracers involved in the correlation function. The equations above assume even multipole index and wrapping of the binned pair counts to .
The difference between the infinite and large finite numbers of bins is not a major accuracy issue. However, the previous methodology for Legendre moments had other practical disadvantages: it required a separate jackknife computation for bins just to tune the shot-noise rescaling, and a fit for the survey correction function — the ratio of pair counts in a real survey and an idealized one (similar to a periodic box with the same volume) – for which a piecewise-polynomial form was used somewhat arbitrarily [29]. We have seen this mismatch as an opportunity to streamline the covariance matrix computation procedure for extensive usage with DESI.
@@
Since the projection in Eq. 3.1 is linear, the covariance matrix for these Legendre moments estimators can be obtained from the -binned one given by Eq. A.1:
@@
A major technical result of this paper is to present a methodology to accumulate this covariance matrix of the Legendre multipoles directly within the summation over point configurations, rather than having to compute and then project the much larger covariance matrix of fine angular bins. For this, several quantities need to be inserted into the sums of Eq. A.2, and we obtain the following 4, 3, and 2-point terms:
@@
Like in Eq. A.2, we include the non-Gaussian higher-point function but null them in the current implementation, which is marked by crossing them out. Sums like practically mean finding the angular bin to which the value belongs and then evaluating the rest only for that one bin. Within the code, we sample a quad, triple, or pair of particles and then accumulate its contribution to all the Legendre multipole moments in its radial bin.
@@
These 4, 3, and 2-point terms can be combined to the full theoretical estimate via
The next sections of this paper only deal with single-tracer covariances, but we provide them since they have been used for cross-correlation analysis [35].
We have found it challenging to design a reasonable and convenient weighting for the jackknife estimates of the Legendre multipole moments. Instead, we have decided to project the -binned jackknife covariance matrix estimate (Eq. (2.7)) similarly to the full covariance (Eq. (3.3)):
and project the theoretical predictions (Eq. (A.3)) in the same fashion, for the just comparison:
with
@@
Similarly, sums of the form practically mean finding the angular bins to which the values belong correspondingly and then evaluating the rest only for that pair of bins. In practice, in the 4-point term, we also omit the disconnected part (), since it complicates the computation and has been found very small in practice.
The key advantage of this new mode is making the Legendre covariance with shot-noise rescaling tuned on jackknives in one go, instead of having to perform two separate runs previously. Additionally, one does not need to fit the survey correction function with a somewhat arbitrary form in non-trivial geometry [29]; instead, the random-random counts are used directly.
4 Validation setup
In this section, we describe how we validate these methods and the RascalC code on applications for DESI. To do so, we apply the data pipeline888The scripts are available at https://github.com/misharash/RascalC-scripts/tree/DESI2024/DESI/Y1, pre and post directories for single-tracer runs. (Fig. 1) to single mock catalogs999These scripts are available at https://github.com/misharash/RascalC-scripts/tree/DESI2024/DESI/Y1/EZmocks/single, pre and post folders., as was done in previous works. This implies using the random files, full and jackknife correlation function estimates specific to that catalog. Similarly to [31], we use effective Zel’dovich (EZ) mocks101010https://github.com/cheng-zhao/EZmock [13, 14], apply compact covariance matrix similarity measures to each of the single-mock runs before and after standard BAO reconstruction (pre- and post-recon) in the observable space (i.e., in terms of correlation function multipoles), and then also project the covariances into the model parameter space (as described in Section 2.3). Earlier papers [27, 28, 29] did not use the full set of the compact similarity measures, did not provide both their expectation values and standard deviations for the perfect theoretical covariance case (Section 2.2) and did not project to parameter space(s). Relative to [31], the new features of this work are:
including quadrupole and hexadecapole moments of the correlation function with the new method (Section 3);
including more tracers (galaxy types): not only luminous red galaxies (LRG) [45] ( bin in this work), but also emission line galaxies (ELG) [46] and magnitude-limited bright galaxy survey (BGS) [47] , corresponding to LRG3, ELG2 and BGS samples in the main BAO paper [7] respectively, as summarized in Table 1;
comparing different separation ranges and sets of multipoles in the observable space (not only the widest one, Section 5.3);
considering more models for Fisher projections: instead of only 1D (isotropic) BAO, now we have 2D BAO, ShapeFit and direct fit (Section 5.4);
as for the data, the covariances have been run separately for North and South galactic caps (NGC and SGC) and then combined as a weighted average assuming those regions are uncorrelated (Appendix B).
@@ 4 Validation setup
To set up the covariances (Fig. 1), we have used 10 mocks for each of the 3 tracers, pre- and post-recon, and 6 different cuts or projections. The mock sample covariance was computed with all 1000 EZmocks in each case. We have decided to only present the mean and standard deviation of the comparison measures across the 10 mocks, the full set of numbers will be available in the supplementary material.
We use the suite of EZmocks calibrated to the clustering of the DESI One-percent survey data [4], cut to the DESI DR1 footprint [15] and run through fast fiber assignment [43].
Effects of fiber assignment may pose additional challenges to the method because it involves anisotropic pair-wise sampling, depending both on the density of the targets and the number of survey passes in the region [44]. We provide the RascalC code with the random catalog and clustering estimate affected by fiber assignment (see the flowchart in Fig. 1), but the expansion leading to Eq. A.2 (and Eq. A.3) uses the survey-wide correlation function(s) to calculate the ensemble averages of products of overdensities, and the shot-noise rescaling (Eq. 2.4) is also global. It is challenging to let them vary without complicating the covariance matrix model too much and introducing too many parameters.
Fiber assignment incompleteness might also cause issues with jackknife. Ideally, one would like each sub-region to have a distribution of the number of passes representative of the full survey. This is challenging to achieve, and the jackknife assignment based on data -means clustering (implemented in pycorr and used in this work) likely does not guarantee that.
We have not endeavored to validate multi-tracer covariances in this work because the snapshot redshifts for most overlapping tracers do not match in the EZmocks suite. The main interest has been in the LRG and ELG overlap in bin [35], but the snapshot redshifts are 1.1 and 0.95 respectively. We expect that the shift of the positions with time will not allow self-consistent cross-correlations. Only part of the QSO come from snapshot redshift 1.1, but that overlap will be even less significant than with ELG.
The reconstruction procedure follows the findings of the DESI DR1 optimal reconstruction task force [33, 34]: the RecSym mode of the IterativeFFTReconstruction algorithm [48] from the pyrecon package111111https://github.com/cosmodesi/pyrecon with smoothing scale of .
@@
The covariance matrices have been created for monopole, quadrupole and hexadecapole in 45 radial bins between 20 and 200 (each 4 wide). We have excluded the bins because they impede the convergence of the covariance matrices, and also because we expect the shot-noise rescaling to become inadequate on small scales.
We also project the covariance matrices into parameter spaces of models according to Section 2.3. We used the following ones implemented in the desilike package121212https://github.com/cosmodesi/desilike:
2D BAO – BAO power spectrum template and polynomial-based broadband modeling (as detailed in [7]), fit to monopole and quadrupole in radial bins spanning ;
Unlike in [31], we exclude the nuisance parameters right away by leaving out the corresponding rows and columns from the covariance matrices given by Eqs. 2.14 and 2.12. The remaining sub-matrix represents the covariance of the parameters of interest marginalized over the others. The RascalC precision matrix is then simply the inverse of this restricted covariance (as the inversion bias is practically negligible for semi-analytical results).
We used a fixed model Jacobian (Eq. 2.13): computed once at the best fit to the mean clustering of all the available mocks (using the mock sample covariance matrix) for each case (tracer, pre- or post-recon, and the model). This is a limitation, which nevertheless allowed us to summarize the comparison measures for all the 10 tested mocks for each setup, and keep the perfect reference simpler.
We also refer the reader to the companion covariance comparison paper [32] for thorough plots of the fit results (not only errorbar estimates, but also parameters’ best values) for each mock catalog with the mock sample covariance against those with RascalC ones. On the flip side, such a detailed approach limited the number of RascalC realizations that could be presented.
5 Results
Having described the methodology for the current RascalC validation on EZmocks in the previous section, we proceed to detail the results of its application.
5.1 Runtime and intrinsic convergence checks
As in [31], we begin by using between the different estimates of each RascalC covariance matrix from separate halves of the Monte-Carlo integration samples for an additional quality assessment before looking at the sample covariance matrices.
We found that the LRG covariance matrices each reached convergence within 4 node-hours on the NERSC Perlmutter supercomputer. The ELG covariance matrices reached convergence within 10 node-hours, while those of BGS reached convergence within 12 node-hours. A single Perlmutter node has a CPU with 128 cores with 256 hyperthreads, all of which are utilized by RascalC. In a few cases, the result was considerably worse, then additional runs were performed, and used either instead or together with the first runs, depending on their .
Relative to [31], these are longer runtimes and yet less converged (the maximum was back then). We believe the key reason is a larger number of observables in the covariance matrix (3 times more – the same number of radial bins, but now 3 multipoles instead of 1). For reference, the comparison of a sample covariance based on 1000 realizations with the true covariance matrix describing the samples’ distribution yields (Table 4). Moreover, the tracers with a higher number density (ELG and especially BGS) are more challenging due to the increasing importance of the 4-point term relative to the 3- and 2-point terms. The more points, the more configurations need to be sampled for good precision of the term.

### v3

@@ 2.2 Comparison measures for covariance matrices
We employ the same three compact measures of covariance matrix similarity as in [35] for detailed comparison and validation with mocks. We invert the final RascalC covariance matrix and not the mock sample covariance matrix because the latter is more noisy and thus less stable to inversion. We also find this ordering more interpretable: some additional properties do not hold with the other one, as noted below. We start by listing the quantities and briefly explaining their meaning.
@@
(2.16) which is used to optimize shot-noise rescaling, and related to the log-likelihood of the sample covariance under the assumption that the RascalC covariance truly describes the distribution of mock clustering measurements121212This log-likelihood relation does not hold if the KL divergence is computed between sample precision (inverse covariance) and theoretical covariance matrices. [28].
@@
(2.17) and are sensitive to deviations in different ways, but their expectation values would not be zero even if RascalC covariance matrices perfectly matched the true underlying covariance describing the distribution of mock realizations.
@@
(2.18) which is distributed like a reduced with degrees of freedom under the assumption that RascalC covariance describes the distribution of the mock clustering measurements131313Again, if we computed the mean reduced between the mock precision and theoretical covariance matrices, it would not follow the reduced distribution exactly.. This measure is more sensitive to the overall scaling between the covariance matrices, whereas deviations in different directions can cancel each other.
It is also important to set expectations for these three comparison measures in the perfect case. By this, we mean comparing the true precision matrix with the sample covariance estimated from multivariate normal samples following the same covariance. The distribution of clustering statistics can be assumed normal. This allows us to test the hypothesis that RascalC can predict the true covariance of the mocks.
We compute the statistical means and standard deviations for our comparison measures in two ways:
Theoretically — using the expressions derived in [35]. It is important to note that the results for are exact while the others are obtained with approximations.
Empirically (Monte-Carlo) — using a large number of multivariate normal samples. We generate chunks of independent samples of -dimensional multivariate normal vectors with a (true) unit covariance matrix141414The value of the true covariance matrix does not matter because it cancels out in each comparison measure (ignoring the numerical instabilities). The dimension, however, is crucially important. Therefore we repeat the procedure for each value of relevant for comparisons in this work.. In each chunk, we estimate the sample covariance and compute the comparison measures between their true (unit) precision matrix and the sample covariance estimate. Finally, we estimate the mean and standard deviation for each comparison measure using the obtained random realizations151515These realizations of covariance matrix comparison measures (for ), as well as all estimates of their means and standard deviations (theoretical, empirical and our fiducial choice between the two for each case), are provided in the supplementary material: doi:10.5281/zenodo.10895161..
We choose a preferred (fiducial) value between the theoretical estimate and the corresponding empirical figure in each case. We prefer the theoretical value unless the difference is more than . Otherwise, we select the empirical estimate. As suspected, we only find deviations for KL divergence (with larger numbers of bins) and (for smaller numbers of bins). The fiducial values are used in Tables 5, 6, 7, 8, 9 and 10 as reference for comparison measures between RascalC precision matrices and sample covariances.
@@ 2.3 Fisher projection to the space of model parameters
Comparing the full covariance matrices for observables (e.g., binned 2PCF or 2PCF multipoles) may be overly generic. In such a case, we consider every arbitrary “direction” in the high-dimensional linear space of these observables. Some of these directions may be unphysical, whereas others can have little or no connection to the data analysis. To highlight a smaller number of meaningful, impactful directions, [52, 53] suggested projecting the covariance matrices into a model-dependent subspace.
We project the covariance matrices to the model parameters through the inverse of the Fisher matrix, following [35] for simplicity. For RascalC results, one can use a simple expression neglecting the inversion bias (Hartlap-like) corrections161616As we mentioned in Section 2.1, such corrections are small for RascalC.:
is the RascalC covariance matrix restricted to the separation and multipole range of the respective model. This range consitutes observables. is the Jacobian of the model, i.e. the matrix of derivatives of radially-binned 2PCF multipoles with respect to the parameter vector :
For the mock sample covariance, we use a similar expression:
is the number of mock samples used to estimate the sample covariance, and is the number of model parameters.
The important difference from Eq. 2.19 is that the sample covariance estimate is more noisy than the semi-analytical model result. Therefore non-linear matrix inversion operations cause significant biases. [54] considered this problem and gave a single (Hartlap-like) correction factor in their Eq. (B6), which we use here. The additional multiplier removes the inversion biases to leading order.
@@ 3 Covariance for projected Legendre moments of 2PCF revisited
Theoretical models of large-scale structure often use multipole (Legendre) moments of the 2-point correlation function instead of its angularly () binned estimates (e.g. [32, 43]). Moreover, the number of multipoles of interest is typically low — monopole, quadrupole, and sometimes hexadecapole [55, 56, 57, 58, 59]171717These references discuss the power spectrum modeling, but the ’th Legendre moment of the correlation function is determined only by the same-order multipole of the power spectrum via a spherical Bessel transform.. It is then convenient to compress the correlation by converting the angularly-binned correlation function (with more than 3 angular bins) to the Legendre moments. Covariance matrices for Legendre multipoles have a lower dimension, which causes fewer numerical problems and makes them easier to estimate directly.
The covariance matrix model for Legendre moments was developed in RascalC previously [31], but this implementation has several disadvantages.
First, it is not directly compatible with jackknives. In practice, producing the Legendre moment covariance with optimal shot-noise rescaling based on data (not relying on a mock sample) requires two separate computations: one for angular () bins with jackknives to tune the shot-noise rescaling, and another to construct the full covariance matrix model for Legendre multipoles. This is an inconvenience when one does not intend to use the angularly-binned correlation function in cosmological inference.
Second, the 2PCF estimation library used in DESI, pycorr181818https://github.com/cosmodesi/pycorr [33], operates under slightly different assumptions. pycorr uses the angularly binned 2-point correlation function (Eq. 2.1) estimates with a large but finite number () of angular bins to compute the radially binned Legendre moments:
where
are the projection factors, which do not depend on radial bins or the tracers involved in the correlation function. The equations above assume even multipole index , and binning in 191919pycorr is capable of binning in , retaining the sign information. But all auto-correlation functions are necessary symmetric, and it is easy to “wrap” the counts to bins in any case as long as the number of bins is even.. In contrast, the RascalC estimators developed earlier assume weighting by Legendre polynomials during pair counting [31]. This is equivalent to using infinitesimally narrow angular bins in Eq. 3.1. The mismatch in assumptions between pycorr and RascalC may not cause significant differences in practice, but it is not desirable.
Third, an additional step was needed to account for realistic survey geometry. The previous RascalC realization relies on the survey correction function — the ratio of pair counts in a real survey and a periodic box with the same volume [60]. This function needed to be modeled for arbitrary angles (). A piecewise-polynomial form has been assumed [31]. Whereas they explain the need for two different polynomials by the particularly strong redshift-space distortions near the line of sight (), the choice of the partition point202020I.e. the boundary between the two polynomials, where they join continuously and smoothly. at has not been motivated and may not be the best. This might be a minor issue as well, but still a source of extra uncertainty.
@@
We have seen an opportunity to address all three issues and streamline the covariance matrix computation procedure for extensive usage with DESI. Since the projection in Eq. 3.1 is linear, the covariance matrix for these Legendre moments estimators can be obtained from the -binned one given by Eq. 2.6:
@@
A major technical result of this paper is a methodology to compute this covariance matrix of the Legendre multipoles directly at the level of the summation over point configurations, rather than having to compute and then project the much larger covariance matrix of fine angular bins. For this, several quantities need to be inserted into the sums of Eq. 2.7, and we obtain the following 4, 3, and 2-point terms:
@@
As in Eq. 2.7, we include the non-Gaussian higher-point functions in these theoretical equations. However, we drop them in the current implementation, which we signify by crossing them out. Sums like practically mean finding the angular bin to which the value belongs and then evaluating the rest only for that one bin. Within the code, we sample a quadruplet, triplet, or pair of particles and then accumulate its contribution to all the Legendre multipole moments in its radial bin.
@@
These 4, 3, and 2-point terms can be combined to the full theoretical estimate analogously to Eq. 2.6, i.e.
This is the single-tracer expression; the version for multiple tracers is provided in Section A.2.
For simplicity, we have decided to reuse the binned jackknife covariance matrix estimate (Eq. (2.10)). An alternative could be a tedious re-derivation of the theoretical jackknife covariance model with some weights for individual jackknife multipole estimators. The method only requires this step to calibrate the shot-noise rescaling parameter. The precise choice of the reference jackknife covariance should not matter as long as the data and model estimators are treated consistently. Consequently, we project the angularly-binned data jackknife covariance matrix (Eq. (2.10)) similarly to the full covariance (Eq. (3.3)):
and do the same with the theoretical prediction (Eq. 2.11):
The jackknife -point terms (Eq. 2.12) accordingly are transformed to
@@
Similarly to Eq. 3.4, sums of the form practically mean finding the angular bins to which the values belong correspondingly and then evaluating the rest () only for that pair of bins. A given set of cells contributes to covariance of one pair of angular bins ( in Eq. 2.7), but all multipoles ().
We have also omitted the small disconnected part of the 4-point jackknife term () in the main code implementation for practical reasons. This part is estimated in bins (Eq. 2.12) by separating it into a product of two sums that need to be computed for each jackknife region [30]. The evaluation becomes less convenient in Legendre multipoles due to the projection factors inserted into the sum over cells/particles. We have computed the disconnected term in a couple of realistic setups by using the technique for bins212121Storing more data due to keeping many (100) angular () bins for each of 60 jackknife regions. The number of angular bins could be reduced, but at the price of potential loss of precision. and projecting the resulting covariance matrix part into the multipole moments (analogously to Eq. 3.6). We have found that the inclusion of the disconnected term does not change the shot-noise rescaling values up to the 6th digit after the decimal point. The optimal shot-noise rescaling parameter is the only link between the disconnected jackknife term and the final covariance matrix. Therefore we concluded that the impact of the disconnected term is practically negligible.
In addition, we give a theoretical justification for neglecting the disconnected term, although not completely strict. The disconnected jackknife term vanishes exactly if either the correlation function is constant, the jackknife regions are identical, or the jackknife counts in each region are the same in each fine bin [30]. An arbitrary survey would not meet any of these conditions exactly, but they likely hold approximately. Therefore the disconnected term is expected to be small.
This concludes the description of the new covariance estimators we apply to DESI data and mocks. To reiterate, the key practical advantage is that a single computation provides the Legendre covariance with shot-noise rescaling tuned on jackknives. Additionally, the new estimators use the random-random counts directly, instead of relying on a survey correction function fit for realistic survey geometry [31]. Therefore we use this method for covariance matrix estimation in the rest of this paper.
4 DESI simulations and methods
In this section, we briefly describe the simulations we use and their processing steps.
4.1 Mocks
In this work, we mainly rely on effective Zel’dovich mocks (EZmocks) [13, 14]. We use the suite of 1000 catalogs representative of DESI DR1 [15]. These mocks are more approximate than those relying on full -body simulations. However, they are fast enough to make a large number of simulations in boxes covering the whole volume of DESI DR1 without replications.
For shot-noise rescaling investigation (Section 6.2) we also used more realistic Abacus-2 mocks. They are based on the AbacusSummit suite of -body simulations [61] produced with the Abacus code [62]. The halos have been identified with the CompaSO halo finder [63]. The galaxy catalogs have been produced within halo occupancy distribution (HOD) formalism using the AbacusHOD framework [64]. More details on the suite representing DESI DR1 can be found in [7]. These simulations have two disadvantages: only 25 realizations and smaller () boxes requiring replications to cover the DESI DR1 volume.
The factors of number and volume are crucial for covariance matrices. The number of mocks sets the relative precision of the sample covariance estimate, which is the only reference we have for comparison. By replications, we mean different parts of the final galaxy catalog made from the same part of the original simulation. These parts can become too strongly correlated and therefore bias the sample covariance estimate. This motivates our choice of EZmocks for the majority of this work.
4.2 Fiber assignment modeling
An important novel aspect of this work is the application of semi-analytical covariance matrices to mocks that incorporate a model of fiber assignment effects.
We mainly use the approximate algorithm called “fast fiber assign” or “fast fiber assignment” (FFA) [48, 46]. It emulates the DESI fiber assignment algorithm using less computational resources [7].
For shot-noise rescaling investigation (Section 6.2) we also used the mock realizations of the DESI fiber assignment algorithm. They involve alternate merged target ledgers, thus the shortcut AMTL [47]. However, this method is prohibitively slow to process all 1000 EZmocks [7].
These fiber assignment algorithms vary in how they assign weights to galaxies and random points to mitigate incompleteness effects. As discussed in section 5.1 of [45], the DESI assignment completeness can be split into two components, one of which can be modeled by assigning weights to randoms. However, the FFA algorithm only determines the total assignment completeness per galaxy. Consequently, the weights per galaxy resulting from FFA have greater variance than for AMTL mocks and the DESI DR1 LSS catalogs. As we mentioned earlier, different weighting methods for data and random points can affect the shot-noise rescaling values (Eq. 2.4).
Effects of fiber assignment may pose additional challenges to the method because it involves anisotropic pair-wise sampling, depending both on the density of the targets and the number of survey passes in the region [46]. We provide the RascalC code with the random catalog and clustering estimate affected by fiber assignment (see the flowchart in Fig. 1), but the expansion leading to Eq. 2.7 (and Eq. 2.12) uses the survey-wide correlation function(s) to calculate the ensemble averages of products of overdensities, and the shot-noise rescaling (Eq. 2.4) is also global. It is challenging to let the correlation function or shot-noise rescaling vary across the survey (with on-sky position or redshift) without complicating the covariance matrix model too much and introducing too many parameters.
Fiber assignment incompleteness might also cause issues with the jackknife. Ideally, we would like each sub-region to have a distribution of the number of passes representative of the full survey. This is challenging to achieve, and the current jackknife assignment does not guarantee that.
4.3 Reconstruction
We apply the semi-analytical covariance matrix estimators before and after standard BAO reconstruction (pre- and post-recon). The reconstruction methodology follows the findings of the DESI DR1 optimal reconstruction task force [37, 38]: the RecSym mode of the IterativeFFTReconstruction algorithm [65] from the pyrecon package222222https://github.com/cosmodesi/pyrecon with smoothing scale of .
4.4 Theoretical modeling and fitting
Our first analysis of interest is DESI 2024 baryon acoustic oscillations [7]. We use the same anisotropic (2D) model with BAO power spectrum template and spline-based broadband terms [43]. The fit uses monopole and quadrupole in radial bins spanning .
The second analysis of interest is DESI 2024 full-shape [9]. This presents more difficulties because it primarily uses power spectra and the methodology has not been standardized for correlation functions. Nevertheless, we use similar models relying on the velocileptors Lagrangian perturbation theory model [66, 67, 56] with maximum freedom and standard prior basis. Two variants correspond to the power spectrum template choice. One approach is “ShapeFit”. It is a compression method using parametric variations of a single power spectrum template evaluated at the reference cosmology [68]. The other is “full modeling”, meaning a linear power spectrum from CLASS Boltzmann code [69]. The fit uses monopole, quadrupole and hexadecapole in radial bins spanning .
These analyses inform our work in two ways.
First, the two fit ranges set our choice of sub-matrices of correlation function multipoles covariances for comparison. In other words, we restrict the covariance matrices to the multipoles and radial bins used in the fitting. This gives us an informed choice for smaller sub-matrices to see which parts of the covariance matrix are captured better by the semi-analytic method.
Second, we project the covariance matrices into parameter spaces of these three models. We do this with the inverse Fisher matrix as described in Section 2.3. It is very important to see that the errorbars on physical parameters are predicted reliably through the RascalC covariances, although this analysis may not generalize to alternative models.
We use the desilike package232323https://github.com/cosmodesi/desilike for the evaluation of all three models (including polynomial emulators for the full-shape models) and fitting the data.
5 Setup for semi-analytical covariance matrices
@@
We consider three galaxy types (tracers) in three different redshift ranges. We use luminous red galaxies (LRG) with , emission line galaxies (ELG) [70] with and magnitude-limited bright galaxy survey (BGS) [71] with . These correspond to LRG3, ELG2 and BGS samples in the main DESI BAO papers [7, 10] respectively. The summary of tracers and redshift bins is provided in Table 1.
We apply the data pipeline242424The data scripts are available at https://github.com/cosmodesi/RascalC-scripts/tree/DESI2024/DESI/Y1, pre and post directories for single-tracer covariances before and after reconstruction respectively. Our analogous single-mock scripts (with only minor differences) are in https://github.com/cosmodesi/RascalC-scripts/tree/DESI2024/DESI/Y1/EZmocks/single, pre and post folders. (Fig. 1) to single mock catalogs, as was done in previous works. This implies using the random files, full and jackknife correlation function estimates specific to that catalog. We use 10252525The number is a compromise between getting good statistics and saving computing time. different realizations of EZmocks to quantify the fluctuations in the semi-analytical results due to realistic random variations in the input quantities.
We process the North and South Galactic Caps (NGC and SGC) separately. DESI DR1 data has been processed in the same manner. This allows different shot-noise rescaling values, reflecting different completeness patterns in these parts of DESI DR1. Then we combine the two covariances into a single matrix for the full survey assuming the Galactic Caps are uncorrelated262626In the EZmocks NGC and SGC are constructed from separate realizations [15], so they are independent. As a result, we do not validate this assumption for data. (Appendix B).
@@
We create the covariance matrices for monopole, quadrupole and hexadecapole in 45 radial bins between 20 and 200 (each 4 wide). We exclude the bins because they impede the convergence of the covariance matrices. Moreover, we expect the shot-noise rescaling to become inadequate on small scales.
6 Intrinsic tests of the method
Having described the setup for the current RascalC application to EZmocks, we detail the quality checks we perform before comparing the results with the mock sample covariance.
6.1 Intrinsic convergence and computation time
We have two specific convergence criteria for each RascalC computation. The first is the eigenvalue test performed by the code on both full and jackknife covariance matrix terms (Eqs. 3.4 and 3.8). The desired condition is that the minimal eigenvalue of the 4-point term is larger than minus the minimal eigenvalue of the 2-point term [30]. The second is the positive definiteness of the final covariance matrix estimate. We do not use the covariance matrix model that fails either of these tests.
In addition, we have a quantitative measure of convergence without strict thresholds. As in [35], we use the comparison measure (Eq. 2.17) between the different estimates of each RascalC covariance matrix from separate halves of the Monte-Carlo integration samples. We found several high outliers compared to other mock realizations of the same tracer.
We repeat the computation on the mocks with the convergence issues mentioned before. Then we post-process the data from the second computation and the data combined from the two computations. We choose the one that first satisfies the two strict criteria and then gives a lower value.
After this, we find that LRG covariance matrices reached convergence within 4 node-hours (512 core-hours272727Hereafter the figure is for the single computation, i.e. effectively twice longer in a few cases. We count physical cores and not hyperthreads. Later profiling also showed a possibility of improvement by using 64 threads and physical cores instead of 128.) on the NERSC Perlmutter supercomputer. The ELG covariance matrices reached convergence within 10 node-hours (1280 core-hours). BGS reached convergence within 12 node-hours (1536 core-hours).
The computations have become longer relative to [35] whereas convergence figures have worsened (i.e. increased). In this previous work the maximum was . The number of 4-point configurations contributing to the sums (Eqs. 3.4 and 3.8) for LRG NGC or SGC in this work () is very close to the analogous number for full LRG in [35] (), where the sums (Eqs. 2.7 and 2.12) were evaluated for a single angular bin.
The processing of each 4-point configuration becomes longer in Legendre mode. This can be expected because, as mentioned earlier, a given configuration contributes only to one pair of angular bins, but to all Legendre multipoles. Other factors, like CPU differences, parallelism efficiency, and a larger number of 3- and 2-point configurations in this work, may also affect the runtimes.
The increase of the convergence measure can be primarily attributed to three times more observables in the covariance matrix. In this work, we have the same number of radial bins as in [35], but three multipoles instead of only monopole. A higher-dimensional covariance matrix requires more configurations sampled for the same relative precision. Additionally, ELG and especially BGS are more challenging due to the increasing importance of the 4-point term relative to the 3- and 2-point terms. The more points, the more configurations need to be sampled for the same precision of the term. Still, the we obtained are significantly smaller than the expected deviation of the mock sample covariance estimate from the true covariance matrix (282828As we show later in Table 5, this number is based on 1000 mock realizations. For a fixed number of samples, the relative precision of the sample covariance also worsens with the number of observables.).